# Topological Photon Transport in Programmable Photonic Processors via Discretized Evolution of Synthetic Magnetic Fields

### v2

@@ 3.1. Minimal geometry for synthetic magnetic fields
Fig. 3 presents snapshots of the population dynamics for a system initialized with unit population at site 1. The evolution spans one complete modulation period $T$. Under CW modulation (Fig. 3a), the evolution according to Eq. (3) drives coherent population transfer along the path $1\rightarrow 2\rightarrow 3\rightarrow 1$. The system operates with Rabi $\pi$-pulses ($JT/3=\pi/2$) to ensure complete transfer between coupled sites. Simulation results reveal near-perfect sequential transfer at the expected times. Experimental measurements yield fidelities of $9.05\pm 0.04\%, 99$$9.09\pm 0.04\%$, and $99.70\pm 0.01\%$.
CCW modulation (Fig. 3b) implements the reversed evolution according to Eq. (3), driving population along $1\rightarrow 3\rightarrow 2\rightarrow 1$. Both simulation and experiment demonstrate comparable efficiency, with fidelities of $99.57\pm 0.02\%$, $98.47\pm 0.06\%$, and $99.78\pm 0.01\%$. The opposite circulation patterns generate artificial magnetic fluxes of opposite signs, directly confirming time-reversal symmetry breaking in our programmable platform.
We also note that the observed circulation depends on the choice of initialization site. Starting from site 1 produces one-period CW/CCW cycles as demonstrated. If instead the system is prepared on site 2 or 3, the population moves only every other substep, so the effective circulation appears reversed and closes only after two periods.
@@ 3.2. Flux-steered interferometry in a two-plaquette lattice
The next step extends the single triangle to a pair of triangular plaquettes that share an edge, forming a two-path interferometer configuration. Both plaquettes are driven in parallel by the same CW three-step sequence $U_{1},U_{2},U_{3}$ with period $T=3\pi/(2J)$. Static phases are programmed on the outer bonds in the form of Eq. (8), which sets loop phases $\Phi_{A}$ and $\Phi_{B}$ around plaquettes A and B. Because the modulation-induced phase $\Phi_{\text{mod}}$ from the CW drive is common to both loops, the observable that steers interference at the shared vertices is the synthetic flux difference:
$\Delta\Phi\equiv\Phi_{A}-\Phi_{B},$ (15)
consistent with the additive structure in Eq. (9) where the common modulation contribution cancels in the difference.
Fig. 4 presents the implementation and readout of this interferometer. Panel a shows the lattice schematic (sites 1-4) where bonds are color-coded to the three substeps of the CW drive ($U_{1}$ red, $U_{2}$ green on the shared edge, and $U_{3}$ blue). The circular arrows label the net loop phases $\Phi_{A}$ and $\Phi_{B}$ around the left and right triangles; individual bond phases are not depicted. The input is prepared in the symmetric superposition $(|1\rangle+|3\rangle)/\sqrt{2}$ so that amplitudes launch into both loops. During one drive period $T$, the sequence in Eq. (3) routes these amplitudes to the shared sites $2$ and $4$ where they interfere. With this initial state, reversing the drive order yields indistinguishable populations at the shared sites, so the protocol’s time-reversal symmetry breaking is obscured in this configuration. Using a different input, such as adding a relative phase between $|1\rangle$ and $|3\rangle$, would reveal the driving direction. Here we focus on the interferometric response.
The resulting interference is shown in panel b as the simulated mid-period populations $P_{2}(T/2)$ and $P_{4}(T/2)$ as $\Delta\Phi$ is swept. The two traces are complementary and cross at $\Delta\Phi=0$, where the populations are balanced. This follows from a two-path picture. During $U_{2}$ the fields at the shared sites are coherent sums of the amplitudes from the two plaquettes that differ only by the programmed flux difference. With the phase origin chosen so that $\Delta\Phi=0$ yields equal populations, the dependence is consistent with:
$P_{2,4}(T/2)=\frac{1}{2}[1\pm V\sin{(\Delta\Phi/2)}],$ (16)
where $0\leq V\leq 1$ is the visibility, the upper sign corresponds to site 2, and the lower to site 4. This form explains the smooth S-shaped exchange of population and the complementarity $P_{2}+P_{4}=1$.
@@ 3.3. Directional transport in a multi-plaquette hexagonal lattice
We now scale to a geometry where intertwined loops come together in a seven-site hexagon surrounding a central site. In the three-site and four-site lattices each step of the drive couples only two sites, so the evolution follows simple two-level dynamics with complete transfer via Rabi $\pi$-pulses. This clear condition no longer applies in lattices where a driving step may couple more than two sites, which breaks the isolated two-level picture and makes the effective Floquet Hamiltonian long-ranged. We therefore identify stable operating points by examining the quasi-energy spectrum of the Floquet operator $U_{F}$. Its eigenvalues take the form $\lambda_{k}=e^{-i\varepsilon_{k}T}$, where the quasi-energies satisfy $0\leq\varepsilon_{k}<2\pi/T$. The stability of the dynamics is set by separation between the quasi-energies. We define the minimal splitting:
$\Delta\varepsilon_{\text{min}}=\min_{i\neq j}|\varepsilon_{i}-\varepsilon_{j}|,$ (17)
which quantifies the smallest gap in the spectrum (distance measured mod $2\pi/T$ on a circle).
In any physical implementation the ideal operator $U_{F}(T)$ is subject to small, random perturbations in control parameters (e.g. heater currents), leading to an implemented operator:
$U_{\text{actual}}(T)=U_{F}(T)+\delta U,$ (18)
where $\delta U$ is a random matrix representing the integrated effect of experimental imperfections over one period. A small gap ($\Delta\varepsilon_{\text{min}}\rightarrow 0$) indicates a near-degeneracy where the system’s final state becomes highly sensitive to $\delta U$, causing the states to mix unpredictably. A larger gap corresponds to a more stable, well-separated evolution, which ensures the evolution is robust against such perturbations. Our strategy is therefore to find the optimal period, $T_{\text{opt}}$, that maximizes this gap:
@@
Fig. 5: Floquet-period optimization and directional transport in a seven-site hexagon. Top-right: schematic of the lattice with color-coded bonds (red/green/blue for $U_{1}/U_{2}/U_{3}$). The system is initialized in state $|1\rangle$. a, Minimal quasi-energy splitting $\Delta\varepsilon_{\text{min}}$ of the three-site Floquet operator $U_{F}(T)$ versus drive period $T$ (dimensionless; $J=\hbar=1$). Peaks at $\tau$, $3\tau$, $5\tau$ mark the Rabi $\pi$ condition and its odd multiples with the maximum splitting reaching $\Delta\varepsilon_{\text{min}}=0.125$. The dips indicate near-degeneracy of quasi-energies. b, In the seven-site case, operating points are identified at $T_{1}=6.7574$, $T_{2}=13.6647$ (used in sim/exp, $\Delta\varepsilon_{\text{min}}=0.119$), and $T_{3}=25.3791$. c,d, CW modulation: simulated and experimental population dynamics over three driving periods ($t=3T$). Oscillatory exchanges are observed within each substep with an overall CW circulation. Violet boxes mark the multi-peak packet and cyan boxes the single-peak packet, with an overlaid arrow indicating the trajectory; these two contributions form a superposition and both advance in the CW sense. e,f, For CCW modulation, the circulation reverses: the same guides show the multi-peak and single-peak packets advancing CCW. g, First harmonic Fourier phase $\phi^{(1)}/2\pi$, defined in Eq. (21), from the measured and simulated dynamics. Solid (sim) and dashed (exp) are overlaid. Gray guide lines with slopes $\nu_{1}^{\text{CCW}}=0.276$ and $\nu_{1}^{\text{CW}}=-0.222$ indicate the extracted per-period winding numbers. Traces show opposite overall trends for CW (red) versus CCW (blue), consistent with broken time-reversal symmetry. Movies of evolutions are provided in the Supplementary Materials.
@@
The seven-site hexagonal lattice presents qualitatively different behavior from our minimal configurations. Fig. 5(c-f) shows population dynamics over three modulation periods $t=3T$, demonstrating oscillatory exchange between coupled sites superimposed on global directional flow. These oscillations migrate around the hexagon with each Hamiltonian switch, producing a clear rotating pattern that follows the modulation direction throughout the evolution. The drive produces sequential site occupations with high population transfer at various stages, while the central site (7) acts mainly as a transient hub and remains only weakly populated overall. The clockwise case is shown in Fig. 5(c,d) and the counterclockwise case in Fig. 5(e,f), where reversing the drive order flips the direction of the rotating pattern. Extended $9T$ dynamics showing more than one full rotation are provided in Supplementary Sec. S4.
As a complementary view of the transport, Fig. 5g shows the first-harmonic Fourier phase extracted from the measured and simulated populations. Assign each site an azimuthal angle $\theta_{k}$ according to the lattice geometry (with the central site taken as $\theta_{7}=0$), and the complex first-harmonic is formed as:
$C^{(1)}(t)=\sum_{k=1}^{7}P_{k}(t)\,e^{i\theta_{k}},$ (20)
where $P_{k}(t)$ is the population at site $k$. Its phase:
$\phi^{(1)}(t)=\arg\!\big{[}C^{(1)}(t)\big{]},$ (21)
provides a measure of the net angular drift and is expected to have an overall linear growth. Expressed in cycles (i.e., $\phi^{(1)}/2\pi$), the phase drifts nearly linearly over the periods; the slope’s sign sets the transport direction, and small kinks appear when population transiently occupies the central site, pulling $C^{(1)}(t)$ toward $\theta=0$. We quantify this drift by an empirical per-period winding number, $\nu_{1}\equiv\frac{1}{2\pi}\frac{d\phi^{(1)}}{d(t/T)}$, i.e., the slope of $\phi^{(1)}/2\pi$ versus $t/T$. From Fig. 5g we obtain $\nu_{1}^{\text{CW}}=-0.222$ and $\nu_{1}^{\text{CCW}}=+0.276$, with the sign reversing under drive inversion. The experimental traces follow closely (correlation $r=0.989$ for CW; $r=0.985$ for CCW). While the seven-site, open system does not enforce quantization, the nonzero and sign-robust $\nu_{1}$ is consistent with a nontrivial per-period phase winding and the observed chiral transport.
@@ 4. Conclusion
In this work, we have demonstrated that discretized Floquet evolution on a programmable photonic processor provides a powerful and highly flexible method for generating synthetic gauge fields. By systematically breaking time-reversal symmetry with time-ordered coupling sequences, we validated the core principles of this approach, progressing from the demonstration of fundamental chiral transport in a single plaquette to coherent, flux-controlled routing in an interferometric geometry, and finally to directional transport in a complex multi-loop lattice. This technique moves beyond static, fabrication-defined implementations by offering a fully programmable route to engineer complex Hamiltonians and explore a wide range of topological phenomena on a single chip.
Supplementary Materials
S1. MZI Transfer Matrix Derivation
The transfer matrix for the Mach-Zehnder interferometer (MZI) used in this work is derived here. The MZI consists of four components in series: an external phase shifter $\varphi$, a first 3-dB ideal coupler, an internal phase shifter $\theta$, and a second 3-dB ideal coupler (Fig. S1).
Fig. S1: Schematic of a Mach-Zehnder interferometer (MZI) composed of two 3-dB MMIs and two thermo-optic phase shifters ($\varphi,\theta$). Inputs and outputs labeled in 1–2, out 1–2.
The transfer matrices for the phase shifters and the ideal 3-dB coupler are defined as:
$T_{\varphi}=\begin{bmatrix}e^{i\varphi}&0\\ 0&1\end{bmatrix},\quad T_{\theta}=\begin{bmatrix}e^{i\theta}&0\\ 0&1\end{bmatrix},\quad T_{\text{C}}=\frac{1}{\sqrt{2}}\begin{bmatrix}1&i\\ i&1\end{bmatrix}$. (S1)
The total transfer matrix is the product of these individual components:
$T_{\text{MZI}}$ $=T_{\text{C}}T_{\theta}T_{\text{C}}T_{\varphi}$ (S2)
[…]
$=\frac{1}{2}\begin{bmatrix}e^{i\varphi}(e^{i\theta}-1)&i(e^{i\theta}+1)\\ ie^{i\varphi}(e^{i\theta}+1)&1-e^{i\theta}\end{bmatrix}$. (S3)
Using the identities $e^{i\theta}-1=2i\sin(\theta/2)e^{i\theta/2}$ and $e^{i\theta}+1=2\cos(\theta/2)e^{i\theta/2}$, leads to the final result:
$T_{\text{MZI}}=ie^{i\theta/2}\begin{bmatrix}e^{i\varphi}\sin(\theta/2)&\cos(\theta/2)\\ e^{i\varphi}\cos(\theta/2)&-\sin(\theta/2)\end{bmatrix}$. (S4)
S2. Phase Shifter Calibration
Accurate realization of the unitary transformations requires calibrating the thermo-optic phase shifters in each MZI. The total phase accumulated in each MZI consists of both an electrically-controlled contribution and a fabrication-induced static offset:
$\alpha_{\text{total}}(I)=\alpha_{\text{elec}}(I)+\alpha_{\text{static}},$ (S5)
where $\alpha$ represents either $\theta$ or $\varphi$ and $I$ is the applied current.
We empirically find that the heater voltage follows $V=aI^{3}+cI+d$, indicating a temperature-dependent resistance $R(I)=aI^{2}+c$ arising from the resistivity increase with heating. The dissipated power $P=I^{2}R(I)$ induces a phase shift through:
$\alpha_{\text{total}}(I)=\beta(aI^{4}+cI^{2})+\alpha_{\text{static}},$ (S6)
where $\beta$ is the thermo-optic efficiency coefficient.
We characterize each MZI through two sequential measurements. First, we perform resistance characterization with current-voltage measurements from 0 to 1.35 mA in 10 steps, extracting coefficients $a$, $c$, and $d$. Second, we route light exclusively through the target MZI and measure transmitted optical power while sweeping the current in 50 steps. Using the resistance model from the first measurement, we calculate the heating power at each current and space these measurements uniformly in power. The optical power follows a sinusoidal modulation with heating power:
$P_{\text{out}}=P_{\text{avg}}[1\pm\cos(\omega P_{\text{heat}}+\phi_{0})],$ (S7)
where $P_{heat}=aI^{4}+cI^{2}$ and the sign $(\pm)$ depends on the measurement configuration, with opposite input-output ports giving the cross configuration (+) and the same ports giving the bar configuration (-). Fitting this response yields $\omega=\beta$ and $\phi_{0}=\alpha_{\text{static}}$, enabling the calculation of the required current for any target phase through numerical solution of Eq. (S6).
This fitting procedure applies to both internal and external phase shifters. While internal phases require routing light exclusively through individual MZIs [31], external phases require extended interferometric configurations for characterization [32].
S3. Photonic chip image
Fig. S2: Photonic chip. Optical microscope image of the programmable photonic processor. The chip measures $16\times 3$ $m$ and integrates 264 elements (phase shifters and MMIs). Deep trenches ($280\times 12$ µm) are etched around the Mach-Zehnder interferometers to reduce thermal crosstalk between neighboring heaters. Inset: magnified view highlighting an MMI, a thermo-optic heater, and surrounding trenches.
S4. Extended simulation of seven-site transport
To extend the main-text results, we simulate the seven-site hexagon over a longer evolution window of $9T$, see Fig. S3. This allows us to confirm that the directional transport is not a short-time artifact but persists across multiple cycles. In this regime, the circulating multi-peak/single-peak structure executes more than one full rotation around the perimeter, demonstrating sustained chiral motion.
Fig. S3: Extended seven-site dynamics. Simulated evolution under CW (left) and CCW (right) drives at the gap-optimized period $T=13.6647$, using 15 steps per period. Horizontal planes mark 0, 3T, 6T, 9T0,\,3T,\,6T,\,9T. The characteristic multi-peak/single-peak pattern described in the main text persists over long times, advancing chirally around the perimeter and completing on the order of two full rotations. Violet boxes mark the multi-peak packet and cyan boxes mark the single-peak packet; for clarity, arrows used in the main-text are omitted here.

### v3

@@ 3.1. Minimal geometry for synthetic magnetic fields
Fig. 3 presents snapshots of the population dynamics for a system initialized with unit population at site 1. The evolution spans one complete modulation period $T$. Under CW modulation (Fig. 3a), the time-ordered sequence defined in Fig. 1a drives coherent population transfer along the path $1\rightarrow 2\rightarrow 3\rightarrow 1$. The system operates with Rabi $\pi$-pulses ($JT/3=\pi/2$) to ensure complete transfer between coupled sites. Simulation results reveal near-perfect sequential transfer at the expected times. Experimental measurements yield target-site populations of $9.05\pm 0.04\%, 99$$9.09\pm 0.04\%$, and $99.70\pm 0.01\%$.
CCW modulation (Fig. 3b) implements the reversed evolution sequence, driving population along $1\rightarrow 3\rightarrow 2\rightarrow 1$. Both simulation and experiment demonstrate comparable efficiency, with measured populations of $99.57\pm 0.02\%$, $98.47\pm 0.06\%$, and $99.78\pm 0.01\%$. The opposite circulation patterns generate artificial magnetic fluxes of opposite signs, directly confirming time-reversal symmetry breaking in our programmable platform.
We also note that the observed circulation depends on the choice of initialization site. Starting from site 1 produces one-period CW/CCW cycles as demonstrated. If instead the system is prepared on site 2 or 3, the population moves only every other substep, so the effective circulation appears reversed and closes only after two periods. Thus, on average the magnetic field is zero, but it depends on both site location and the phase of the drive.
@@ 3.2. Flux-steered interferometry in a two-plaquette lattice
The next step extends the single triangle to a pair of triangular plaquettes that share an edge, forming a two-path interferometer configuration. Both plaquettes are driven by the CW time-ordered sequence with period $T=3\pi/(2J)$. Static phases are programmed on the outer bonds to set loop phases $\Phi_{A}$ and $\Phi_{B}$ around plaquettes A and B. Because the modulation-induced phase $\Phi_{\text{mod}}$ is common to both loops, the relevant quantity that steers interference at the shared vertices is the synthetic flux difference $\Delta\Phi\equiv\Phi_{A}-\Phi_{B}$. This cancellation of the modulation phase allows us to isolate the static contribution.
Fig. 4 presents the implementation and readout of this interferometer. Panel a shows the lattice schematic (sites 1-4) where bonds are color-coded to the three substeps of the CW drive ($U_{1}$ red, $U_{2}$ green on the shared edge, and $U_{3}$ blue). The circular arrows label the net loop phases $\Phi_{A}$ and $\Phi_{B}$ around the left and right triangles; individual bond phases are not depicted. The input is prepared in the symmetric superposition $(|1\rangle+|3\rangle)/\sqrt{2}$ so that amplitudes launch into both loops. During one drive period $T$, the CW sequence routes these amplitudes to the shared sites $2$ and $4$ where they interfere. With this initial state, reversing the drive order yields indistinguishable populations at the shared sites, so the protocol’s time-reversal symmetry breaking is obscured in this configuration. Using a different input, such as adding a relative phase between $|1\rangle$ and $|3\rangle$, would reveal the driving direction. Here we focus on the interferometric response.
The resulting interference is shown in panel b as the simulated mid-period populations $P_{2}(T/2)$ and $P_{4}(T/2)$ as $\Delta\Phi$ is swept. The two traces are complementary and cross at $\Delta\Phi=0$, where the populations are balanced. This follows from a two-path picture. During $U_{2}$ the fields at the shared sites are coherent sums of the amplitudes from the two plaquettes that differ only by the programmed flux difference. With the phase origin chosen so that $\Delta\Phi=0$ yields equal populations, the dependence is consistent with $P_{2,4}(T/2)=\frac{1}{2}[1\pm V\sin{(\Delta\Phi/2)}]$, where $0\leq V\leq 1$ is the visibility, the upper sign corresponds to site 2, and the lower to site 4. This form explains the smooth S-shaped exchange of population and the complementarity $P_{2}+P_{4}=1$.
@@ 3.3. Directional transport in a multi-plaquette hexagonal lattice
We now scale to a geometry where intertwined loops come together in a seven-site hexagon surrounding a central site. In the three-site and four-site lattices each step of the drive couples only two sites, so the evolution follows simple two-level dynamics with complete transfer via Rabi $\pi$-pulses. This clear condition no longer applies in lattices where a driving step may couple more than two sites, which breaks the isolated two-level picture and makes the effective Floquet Hamiltonian long-ranged. We therefore identify stable operating points by examining the quasi-energy spectrum of the Floquet operator $U_{F}$. Its eigenvalues take the form $\lambda_{k}=e^{-i\varepsilon_{k}T}$, where the quasi-energies satisfy $0\leq\varepsilon_{k}<2\pi/T$. The stability of the dynamics is set by the separation between the quasi-energies, defined as the minimal splitting $\Delta\varepsilon_{\text{min}}=\min_{i\neq j}|\varepsilon_{i}-\varepsilon_{j}|$, which quantifies the smallest gap in the spectrum (distance measured mod $2\pi/T$ on a circle).
In any physical implementation the ideal operator $U_{F}(T)$ is subject to small, random perturbations in control parameters (e.g. heater currents), leading to an implemented operator $U_{\text{actual}}(T)=U_{F}(T)+\delta U$, where $\delta U$ is a random matrix representing the integrated effect of experimental imperfections over one period. A small gap ($\Delta\varepsilon_{\text{min}}\rightarrow 0$) indicates a near-degeneracy where the system’s final state becomes highly sensitive to $\delta U$, causing the states to mix unpredictably. A larger gap corresponds to a more stable, well-separated evolution, which ensures the evolution is robust against such perturbations. Our strategy is therefore to find the optimal period, $T_{\text{opt}}$, that maximizes this gap:
@@
Fig. 5: Floquet-period optimization and directional transport in a seven-site hexagon. Top-right: schematic of the lattice with color-coded bonds (red/green/blue for $U_{1}/U_{2}/U_{3}$). The system is initialized in state $|1\rangle$. a, Minimal quasi-energy splitting $\Delta\varepsilon_{\text{min}}$ of the three-site Floquet operator $U_{F}(T)$ versus drive period $T$ (dimensionless; $J=\hbar=1$). Peaks at $\tau$, $3\tau$, $5\tau$ mark the Rabi $\pi$ condition and its odd multiples with the maximum splitting reaching $\Delta\varepsilon_{\text{min}}=0.125$. The dips indicate near-degeneracy of quasi-energies. b, In the seven-site case, operating points are identified at $T_{1}=6.7574$, $T_{2}=13.6647$ (used in sim/exp, $\Delta\varepsilon_{\text{min}}=0.119$), and $T_{3}=25.3791$. c,d, CW modulation: simulated and experimental population dynamics over three driving periods ($t=3T$). Oscillatory exchanges are observed within each substep with an overall CW circulation. Violet boxes mark the multi-peak packet and cyan boxes the single-peak packet, with an overlaid arrow indicating the trajectory; these two contributions form a superposition and both advance in the CW sense. e,f, For CCW modulation, the circulation reverses: the same guides show the multi-peak and single-peak packets advancing CCW. g, First harmonic Fourier phase $\phi^{(1)}/2\pi$ from the measured and simulated dynamics. Solid (sim) and dashed (exp) are overlaid. Gray guide lines with slopes $\nu_{1}^{\text{CCW}}=0.276$ and $\nu_{1}^{\text{CW}}=-0.222$ indicate the extracted per-period winding numbers. Traces show opposite overall trends for CW (red) versus CCW (blue), consistent with broken time-reversal symmetry. Movies of evolutions are provided in the Supplementary Materials.
@@
The seven-site hexagonal lattice presents qualitatively different behavior from our minimal configurations. Fig. 5(c-f) shows population dynamics over three modulation periods $t=3T$, demonstrating oscillatory exchange between coupled sites superimposed on global directional flow. These oscillations migrate around the hexagon with each Hamiltonian switch, producing a clear rotating pattern that follows the modulation direction throughout the evolution. The drive produces sequential site occupations with high population transfer at various stages, while the central site (7) acts mainly as a transient hub and remains only weakly populated overall. The clockwise case is shown in Fig. 5(c,d) and the counterclockwise case in Fig. 5(e,f), where reversing the drive order flips the direction of the rotating pattern. Extended $9T$ dynamics showing more than one full rotation are provided in section S8.
As a complementary view of the transport, Fig. 5g shows the first-harmonic Fourier phase extracted from the measured and simulated populations. Assign each site an azimuthal angle $\theta_{k}$ according to the lattice geometry (with the central site taken as $\theta_{7}=0$), and the complex first-harmonic is formed as $C^{(1)}(t)=\sum_{k=1}^{7}P_{k}(t)\,e^{i\theta_{k}}$, where $P_{k}(t)$ is the population at site $k$. Its phase $\phi^{(1)}(t)=\arg\!\big[C^{(1)}(t)\big]$ provides a measure of the net angular drift and is expected to have an overall linear growth. Expressed in cycles (i.e., $\phi^{(1)}/2\pi$), the phase drifts nearly linearly over the periods; the slope’s sign sets the transport direction, and small kinks appear when population transiently occupies the central site, pulling $C^{(1)}(t)$ toward $\theta=0$. We quantify this drift by an empirical per-period winding number, $\nu_{1}\equiv\frac{1}{2\pi}\frac{d\phi^{(1)}}{d(t/T)}$, i.e., the slope of $\phi^{(1)}/2\pi$ versus $t/T$. From Fig. 5g we obtain $\nu_{1}^{\text{CW}}=-0.222$ and $\nu_{1}^{\text{CCW}}=+0.276$, with the sign reversing under drive inversion. The experimental traces follow closely (correlation $r=0.989$ for CW; $r=0.985$ for CCW). While the seven-site, open system does not enforce quantization, the nonzero and sign-robust $\nu_{1}$ is consistent with a nontrivial per-period phase winding and the observed chiral transport.
@@ 4. Conclusion
In this work, we have demonstrated that discretized Floquet evolution on a programmable photonic processor provides a powerful and highly flexible method for generating synthetic gauge fields. By systematically breaking time-reversal symmetry with time-ordered coupling sequences, we validated the core principles of this approach, progressing from the demonstration of fundamental chiral transport in a single plaquette to coherent, flux-controlled routing in an interferometric geometry, and finally to directional transport in a complex multi-loop lattice. This technique moves beyond static, fabrication-defined implementations by offering a fully programmable route to engineer complex Hamiltonians and explore a wide range of driven phenomena on a single chip.